Linear Algebra. Please do problem 4 parts a-e. I need all steps clear and concise. I will rate you
4. Find all eigenvalues of the following matrices and find one eigenvector associated with each eigenvalue.
(a) A = = [9] (b) A- والی با 2 (c) A (d) A = (e) A = 11 EC Exercise 55 from Section 6.3 on page 378.
Answer #1

A=[i Zhapetostic Poly nomal 1A-17) - 1-) o 01-1 = x -21+1=0 12-1-) H=0 =) +C+1)-1(2-1)=0 .-) 1=1 / Eigeon vecty Asscrialed aEigears) & A = [os] characterstic Polynomial 1A-11) = 3-4 ] 0-3-1 - 3+1-6=0 » ?+31-21-6=0 =) *[++3) -2(173)=0 - +3) (-2)=0 -)= 3+1=0 = 5=) =)) -1,7 Eigen recty when it - =) JA-I) - CILD) 1-17:0 G -2X220 = x2 = 0 +) X, CO E[ 7 = span wahrs [:7-CI+)-13--- Xa [1] Eigern vecter = span {[-]} If d=3 1 8-21) – (7-201=[:] -2, tl=6 =) 2=X2 Eigen rects span & D